OpenStudy (anonymous):

can 3/20 at its lowest term ? please help !! @TheSmartOne

OpenStudy (anonymous):

3/20 will not reduce any further

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

sure thing :)

OpenStudy (anonymous):

how about 12/40 does that have a lowest term ?

OpenStudy (anonymous):

yes...now that will reduce. When you have 2 even numbers you can start by dividing each by 2 giving you : 12/40 = 6/20 (divide again by 2) giving you : 3/10...which can reduce no more. So the simpliest term of 12/40 = 3/10

OpenStudy (anonymous):

Wow you must love math lol . do you mind if I ask you three more questions ?

OpenStudy (anonymous):

ask away...I like some math...algebra is ok, geometry...well, I hate geometry

OpenStudy (anonymous):

lol okay , -6/11*3/4 does that have a lowest form ?

OpenStudy (anonymous):

-6/11 * 3/4 --- when multiplying fractions, just multiply straight across..but pay attention to your signs. When the signs are the same, it is positive and when the signs are different, it is negative. -6 3 -18 -9 --- x --- = ---- which reduces to ---- <== this does not reduce 11 4 44 22

OpenStudy (anonymous):

so lowest term of : -6/11 * 3/4 = -9/22

OpenStudy (anonymous):

I am sorry but time has slipped up on me...I have to go. I will be back on later.

OpenStudy (anonymous):

ok

TheSmartOne (thesmartone):

Sorry I wasn't online when yo tagged me.